# Structurally damped semilinear evolution equation for positive operators on Hilbert space

### v1

@@ Abstract.
In this study, we analyze a semilinear damped evolution equation under different damping conditions, including the undamped $(\theta=0)$, effectively damped $(0<2\theta<\sigma)$, critically damped $(2\theta=\sigma)$, and non-effectively damped $(\sigma<2\theta\leq 2\sigma)$. The analysis is conducted in two parts; the present article is devoted to examining decay estimates of solutions to the linear evolution equation governed by a self-adjoint, positive operator $\mathcal{L}$ with discrete spectrum subject to initial Cauchy data of minimal regularity. Specifically, we consider the Cauchy problem:
@@
in different damping conditions. Furthermore, we demonstrate that the decay rates of the associated solutions improve with the regularity of the initial Cauchy data. As an application of the decay estimates, we also demonstrate the global existence (in time) of the solution in certain cases, taking into account polynomial-type nonlinearity. In the 2nd article, we will address the remaining instances where global existence cannot be assured and instead present findings on local existence and possible blow-up results.
@@ 1. Introduction
Let $\mathcal{L}$ be a densely defined self-adjoint linear operator on a separable Hilbert space $\mathcal{H}$ with the discrete spectrum
@@
A crucial classification introduced in [DE16] distinguishes between effective and non-effective damping. The damping is said to be:
@@
  • critical (scale-invariant) when $2\theta=\sigma$; and
  • non-effective if $2\theta\in(\sigma,2\sigma]$.
@@
This study explores all four cases, analyzing how decay rates and solution regularity differ depending on the damping effectiveness. But particular attention is paid to damped cases to understand how damping impacts the decay estimates. The case of effective damping for decay is thoroughly investigated but mainly in Euclidean settings; see [Tuan2019, DABBICCO2014, Narazaki2013]. For the critical/scale-invariant case, one can refer to [PHAM2015]. Furthermore, the authors in [Ikheta2014, Abbico2022, Tuan2019a] investigate the decay estimates and long-time behavior of the solutions in the case of non-effective damping.
@@
  • Section 2 states the main results, starting with decay estimates for the linear homogeneous problem under minimal regularity assumptions on initial data, then extending to enhanced decay with higher regularity.
@@
  • Section 5 provides the proofs of the decay estimates in various damping regimes. Then we will also highlight the main steps for the case of higher regularity to obtain sharper decay estimates.
@@ 2. Main results
with the bare minimal assumptions of regularity on the initial Cauchy data, which are $u_{0},u_{1}\in\mathcal{H}$. Then, we will demonstrate that the decay rates of the associated solutions improve with the regularity of the initial Cauchy data and additional assumptions on the self-adjoint and positive operator $\mathcal{L}$.
@@
  • The value of $\delta$ varies among theorems based on assumptions or estimates. To make them easier to read, we utilize the same symbol to denote them throughout the text.
With the notational groundwork in place, we are now highlighting the article’s main results. We start our analysis with the undamped case, which serves as a baseline for comparing the different damped regimes detailed later.
@@
As we can observe in the undamped case, a certain degree of Sobolev regularity in the initial data is typically necessary to obtain meaningful decay estimates on the solution. We will see in the following investigation of the damped cases that even when the initial data is of minimal regularity, i.e., $u_{0},u_{1}\in\mathcal{H}$, the presence of damping enables us to deduce at least a degree of decay for solutions. Let’s begin with the case of effective damping:
@@
Let $2\theta\in(\sigma,2\sigma]$, $\beta>0$, and $k\in\mathbb{N}$. Let $u$ be the solution of the following homogeneous Cauchy problem (2.1), then the solution $u$ satisfies the following estimates
@@
Here are some crucial observations from the decay estimates in the decay estimates:
  • In contrast to the Euclidean case, we obtain exponential decay for higher-order time and spatial derivatives of the solution in the damped cases, at least for large timeframes $t\geq 1$. This is because of the discrete spectrum of the positive operator $\mathcal{L}$.
@@
  • In certain cases, we also witness a complete absence of decay.
Thus, the absence of decay or the presence of polynomial-time decay in the interval $t\in(0,1]$ is generally inadequate to demonstrate global well-posedness. In such situations, one frequently imposes further assumptions on the original data—such as enhanced regularity or very small conditions to guarantee the existence of at least a local-in-time solution. After local existence is established, the study naturally turns to long-term behavior, such as whether the solution may be extended worldwide or blows up finitely. In order to expect more decay from the solution, we now permit the initial Cauchy data to be more regular.
Theorem 2.6.
Let $\theta\geq 0,\sigma,\beta>0,$ and $k\in\mathbb{N}$. Let $u$ be the solution of the homogeneous Cauchy problem (2.1). Then
@@
  2. (2)
for some $\delta>0$ and for all $t>0$.
@@
By introducing more regularity to the initial Cauchy data, we improve the decay rate in the aforementioned theorem from polynomial to exponential but it is important to note that we still do not get any decay for $\|u\|_{\mathcal{H}}$
  • in undamped case with Sobolev regular initial Cauchy data; see Theorem 2.1;
  • in damped cases with minimal regularity $u_{0},u_{1}\in\mathcal{H}$; see Theorem 2.3, 2.4, and 2.5; and
  • even in damped cases with additional Sobolev regular initial Cauchy data; see Theorem 2.6.
  • Also, there is no decay in the first time derivative.
The aforementioned scenario contrasts with the Euclidean case, where this kind of behavior is not commonly seen in classical wave equation but the explanation is somewhat technical and arises out of the Plancherel formula which is due to the possibility of vanishing $\lambda_{\xi}$ for some $\xi$; it will become clearer after the proof is presented.
Theorem 2.8.
Let $\theta\geq 0,\sigma,\beta>0,$ and $k\in\mathbb{N}$. Assume that $\mathcal{L}$ be a self-adjoint and strictly positive operator and $u$ be the solution of the homogeneous Cauchy problem (2.1). Then
  1. (1)
  2. (2)
Remark 2.9.
Given the previous explanation, the aforementioned theorem also aligns with the decay estimates found in [mic-wave2].
@@ 3. Preliminaries
The key components of Fourier analysis in relation to our indicated space and the self-adjoint, positive operator $\mathcal{L}$ will be reviewed in this section, see [DMT17, Ruz16, ruzhansky2017nonharmonic, MR&JPVR] for a thorough exposition. More precisely, this section defines the $\mathcal{L}$-Fourier transform, $\mathcal{L}$-convolution and the Sobolev spaces associated with the operator $\mathcal{L}$ and its properties.
@@
$|\phi(\xi)|\leq C_{\phi,N}\lambda_{\xi}^{-N}\quad\text{for all }\xi\in\mathcal{I}.$
The topology on $\mathcal{S}(\mathcal{I})$ is given by the seminorms $p_{k},k\in\mathbb{N}_{0}$ and is defined by
$p_{k}(\phi):=\sup_{\xi\in\mathcal{I}}\lambda_{\xi}^{k}|\phi(\xi)|.$
@@
[…]
@@
The Parseval’s identity and the Plancherel’s formula are given by
$\langle{f_{1}},{f_{2}}\rangle_{\mathcal{H}}=\langle\widehat{f_{1}},\widehat{f_{2}}\rangle_{\ell^{2}({\mathcal{I}})}\quad\text{and}\quad\|f\|_{\mathcal{H}}=\|\widehat{f}\|_{\ell^{2}({\mathcal{I}})},$ (3.1)
respectively. The $\mathcal{L}$-convolution of $f_{1},f_{2}\in H_{\mathcal{L}}^{\infty}$ is defined by
@@
We now recall Sobolev spaces generated by the operator $\mathcal{L}$. For $f\in\mathcal{D}^{\prime}(\mathcal{L})$ and $s>0,$ we say that $f\in\mathcal{H}_{\mathcal{L}}^{s}$ iff $\lambda_{\xi}^{\frac{s}{2}}\widehat{f}(\xi)\in\ell^{2}({\mathcal{I}})$ and the associated Sobolev norm is defined as
$\|f\|_{\mathcal{H}_{\mathcal{L}}^{s}}:=\left(\sum_{\xi\in\mathcal{I}}\lambda_{\xi}^{s}|\widehat{f}(\xi)|^{2}\right)^{1/2}.$
@@ 4. Preparatory Estimates
[…]
@@
where $*_{\mathcal{L}}$ is the convolution defined in (3.2), equivalently
@@
$\partial^{k}_{t}\widehat{E}_{1}(t,\xi)$ =$=\begin{cases}\frac{(\tau_{+})^{k}e^{\tau_{+}t}-(\tau_{-})^{k}e^{\tau_{-}t}}{\tau_{+}-\tau_{-}}&\text{if }\tau_{+}\neq\tau_{-},\\ (\tau_{+})^{k-1}(k+t\tau_{+})e^{\tau_{+}t}&\text{if }\tau_{+}=\tau_{-}.\end{cases}$ (4.10)
Remark 4.1.
  1. (1)
    For $x\in[0,\frac{1}{4}]$, the following inequalities hold:
    $-4x\leq-1+\sqrt{1-4x}\leq-2x\quad\text{and}\quad-2\leq-1-\sqrt{1-4x}\leq-1.$ (4.11)
  2. (2)
    For any $\gamma,\delta>0$, the function $f(t)=t^{\gamma}e^{-\delta t}$ and $f(t)=(1+t)^{\gamma}e^{-\delta t}$ is bounded on $[0,\infty)$. Furthermore, for any $\gamma>0$ and $\beta>\delta>0$, we have
    for some positive constant $C(\gamma,\delta)>0$.
  3. (3)
    Since ensuring that $\delta$ stays independent of $\xi\in\mathcal{I}$ is adequate for our calculations, we will suppress the dependence of $\delta$ on the parameters $a$ and $\gamma$.
@@ 5. Proof of main results
This section presents the proof of our main results, which is reliant upon the estimates of $\widehat{E}_{0}(t,\xi)$ and $\widehat{E}_{1}(t,\xi)$, as well as their spatial and temporal derivatives. The proofs of Theorem 2.6 and 2.8 will be presented gradually, accompanied by a set of remarks subsequent to the derivation of decay estimates in each respective case.
Note: The constant $\delta$ may change from line to line in this section. However, we don’t continuously relabel it to prevent notational clutter. Only the presence of a positive $\delta$ is necessary for our aims.
@@
where we have utilized the relation (4.11) to estimate the exponential factor and the fact that $\{\lambda_{\xi}\}_{\xi\in\mathcal{I}}$ is discrete, which allows us to bound $\frac{1}{\sqrt{\frac{1}{4}-\lambda_{\xi}^{\sigma}}}$. Consequently, the representation (4.6) along with the above estimates gives
@@
since $\lambda_{\xi}^{\beta}$ is bounded on $\mathcal{R}_{2}$. Recalling the relation (4.10) for $\widehat{E}_{1}(t,\xi)$, we have
@@
for some modified $\delta>0$, where we have utilize the inequalities (4.11), (4.13), and the fact that $\frac{\left(-\frac{1}{2}+\sqrt{\frac{1}{4}-\lambda_{\xi}^{\sigma}}\right)^{k}}{2\sqrt{\frac{1}{4}-\lambda_{\xi}^{\sigma}}}$ and $\frac{\left(-\frac{1}{2}-\sqrt{\frac{1}{4}-\lambda_{\xi}^{\sigma}}\right)^{k}}{2\sqrt{\frac{1}{4}-\lambda_{\xi}^{\sigma}}}$ are bounded on $\mathcal{R}_{2}$. Consequently, using the relation (4.9), we obtain
@@
since $\lambda_{\xi}^{\sigma}$ is bounded on $\mathcal{R}_{2}$. Combining the above estimate with the representation (4.8), we get
@@
$|\lambda_{\xi}^{\beta}\widehat{E}_{0}(t,\xi)|\lesssim\lambda_{\xi}^{\beta}e^{-\frac{1}{2}t}\quad\text{and}\quad|\lambda_{\xi}^{\beta}\widehat{E}_{1}(t,\xi)|\lesssim\frac{\lambda_{\xi}^{\beta}}{\sqrt{\lambda_{\xi}^{\sigma}-\frac{1}{4}}}e^{-\frac{1}{2}t}\lesssim\lambda_{\xi}^{\beta-\frac{\sigma}{2}}e^{-\frac{1}{2}t}.$
Consequently, using the above estimates with (4.6) we obtain
@@
  • The Remark 2.7 is now more understandable. Specifically, estimate (5.15) shows that the regions $\mathcal{R}_{2}$, $\mathcal{R}_{3}$, and $\mathcal{R}_{4}$ experience exponential decay. Nevertheless, the possibility of decay in the total energy norm $\|u(t)\|_{\mathcal{H}}$ is obstructed by the lack of any decay in $\mathcal{R}_{1}$.
@@
$\widehat{E}_{0}(t,\xi)$ =$=\begin{cases}1&\text{if }\xi\in\mathcal{R}_{1},\\ \left[\cosh\left(\sqrt{\frac{1}{4}\lambda_{\xi}^{2\theta}-\lambda_{\xi}^{\sigma}}t\right)+\frac{\lambda_{\xi}^{\theta}\sinh\left(\sqrt{\frac{1}{4}\lambda_{\xi}^{2\theta}-\lambda_{\xi}^{\sigma}}t\right)}{2\sqrt{\frac{1}{4}\lambda_{\xi}^{2\theta}-\lambda_{\xi}^{\sigma}}}\right]e^{-\frac{1}{2}\lambda_{\xi}^{\theta}t}&\text{if }\xi\in\mathcal{R}_{2},\\ \left[1+\frac{1}{2}\lambda_{\xi}^{\theta}t\right]e^{-\frac{1}{2}\lambda_{\xi}^{\theta}t}&\text{if }\xi\in\mathcal{R}_{3},\\ \left[\cos\left(\sqrt{\lambda_{\xi}^{\sigma}-\frac{1}{4}\lambda_{\xi}^{2\theta}}t\right)+\frac{\lambda_{\xi}^{\theta}\sin\left(\sqrt{\lambda_{\xi}^{\sigma}-\frac{1}{4}\lambda_{\xi}^{2\theta}}t\right)}{2\sqrt{\lambda_{\xi}^{\sigma}-\frac{1}{4}\lambda_{\xi}^{2\theta}}}\right]e^{-\frac{1}{2}\lambda_{\xi}^{\theta}t}&\text{if }\xi\in\mathcal{R}_{4},\end{cases}$ (5.18)
@@
$\widehat{E}_{1}(t,\xi)$ =$=\begin{cases}t&\text{if }\xi\in\mathcal{R}_{1},\\ \frac{\sinh\left(\sqrt{\frac{1}{4}\lambda_{\xi}^{2\theta}-\lambda_{\xi}^{\sigma}}t\right)}{\sqrt{\frac{1}{4}\lambda_{\xi}^{2\theta}-\lambda_{\xi}^{\sigma}}}e^{-\frac{1}{2}\lambda_{\xi}^{\theta}t}&\text{if }\xi\in\mathcal{R}_{2},\\ te^{-\frac{1}{2}\lambda_{\xi}^{\theta}t}&\text{if }\xi\in\mathcal{R}_{3},\\ \frac{\sin\left(\sqrt{\lambda_{\xi}^{\sigma}-\frac{1}{4}\lambda_{\xi}^{2\theta}}t\right)}{\sqrt{\lambda_{\xi}^{\sigma}-\frac{1}{4}\lambda_{\xi}^{2\theta}}}e^{-\frac{1}{2}\lambda_{\xi}^{\theta}t}&\text{if }\xi\in\mathcal{R}_{4}.\end{cases}$ (5.19)

### v2

@@ Abstract.
In this study, we analyze the following semilinear damped evolution equation associated with a self-adjoint, positive operator $\mathcal{L}$ on a separable Hilbert space $\mathcal{H}$ with discrete spectrum which, reads as follows:
@@
under different damping conditions, including the undamped $(\theta=0)$, effectively damped $(0<2\theta<\sigma)$, critically damped $(2\theta=\sigma)$, non-effectively damped $(\sigma<2\theta<2\sigma)$, and viscoelastic type damped $(\theta=\sigma)$. This article is devoted to examining the decay estimates of solutions to the above linear evolution equation subject to initial Cauchy data with Sobolev regularity. Furthermore, we also observe loss of decay with relaxation of regularity in some case. As an application of the decay estimates, we also demonstrate the global existence (in time) of the solution in certain cases, taking into account polynomial-type nonlinearity.
@@ 1. Introduction
Let $\mathcal{L}$ be a densely defined self-adjoint, linear, positive operator on a separable Hilbert space $\mathcal{H}$ with the discrete spectrum
@@
A crucial classification introduced in [classif] distinguishes between effective and non-effective damping. The damping is said to be:
@@
  • critical (scale-invariant) when $2\theta=\sigma$;
  • non-effective if $2\theta\in(\sigma,2\sigma)$; and
  • viscoelastic type damping if $\theta=\sigma$.
@@
This study explores all five cases, analyzing how decay rates and solution regularity differ depending on the damping effectiveness. The undamped case has been thoroughly investigated in both Euclidean and non-Euclidean settings; see [GR2, mic-wave2, mic-wave1, palmieri-2] The case of effective damping for decay is thoroughly investigated but mainly in Euclidean settings; see [Tuan2019, DABBICCO2014, Narazaki2013]. For the critical/scale-invariant case, one can refer to [PHAM2015]. Furthermore, the authors in [Ikheta2014, Abbico2022, Tuan2019a] investigate the decay estimates and long-time behavior of the solutions in the case of non-effective damping. The case of viscoelastic type damping is also considered in several settings; see [Shyam:visco, Ikehata, IKEHATA2017228, LIU2021663].
@@
  • Section 2 states the main results, starting with decay estimates for the linear homogeneous problem under Sobolev regularity assumptions on initial data, then we will observe some loss of decay with relaxation in regularity.
@@
  • Section 5 provides the proofs of the main results along with necessary remarks.
@@ 2. Main results
  • $\|(u,v)\|_{\mathcal{D}}:$ For $\mathcal{D}=X\times Y$, where $X$ and $Y$ are normed linear space, we introduce the norm
    $\|(u,v)\|_{\mathcal{D}}:=\|u\|_{X}+\|v\|_{Y}.$
  • The constant $\delta$ below may change from line to line in this paper. However, we don’t continuously relabel it to prevent notational clutter. Only the presence of a positive $\delta$ independent of $\xi\in\mathcal{I}$ is necessary for our aims.
With the notational groundwork in place, we now highlight the main results of the article. We start our analysis with the undamped case, which serves as a baseline for comparing the different damped regimes detailed later.
@@
Remark 2.3.
The estimate $\|u(t)\|_{\mathcal{H}}$ witness a complete absence of decay even in the presence of Sobolev data but the explanation is somewhat technical and arises out of the Plancherel formula which is due to the possibility of vanishing $\lambda_{\xi}$ for some $\xi$; it will become clearer after the proof is presented.
The following theorem provides the decay estimates for effective damping, critical damping, non-effective damping, and viscoelastic type damping
Theorem 2.4.
Let $\theta>0,\sigma,\beta>0,$ and $k\in\mathbb{N}$. Let $u$ be the solution of the homogeneous Cauchy problem (2.1). Then
  1. (1)
  2. (2)
for some $\delta>0$ and for all $t>0$.
Theorem 2.5.
Let $\theta,\beta\geq 0,\sigma>0,$ and $k\in\mathbb{N}$. Assume that $\mathcal{L}$ be a self-adjoint and strictly positive operator and $u$ be the solution of the homogeneous Cauchy problem (2.1). Then
  1. (1)
  2. (2)
Remark 2.6.
The decay estimates in the aforementioned theorem are also aligned with those found in [mic-wave2], given the explanation provided before stating the above theorem.
As we can observe in the Theorem 2.4 and 2.5, a certain degree of Sobolev regularity in the initial data is typically necessary to obtain meaningful decay estimates on the solution. However, in the following theorem, it is observed that even when only the minimal regularity of the initial Cauchy data is utilized, i.e., $u_{0},u_{1}\in\mathcal{H}$, some decay can still be obtained in the cases of effective damping, critical damping, and non-effective damping. Let’s begin with the case of effective damping:
@@
Let $2\theta\in(\sigma,2\sigma)$, $\beta>0$, and $k\in\mathbb{N}$. Let $u$ be the solution of the following homogeneous Cauchy problem (2.1), then the solution $u$ satisfies the following estimates
@@
Remark 2.10.
Here are some crucial observations from the decay estimates:
  1. (1)
    In contrast to the classical wave equation in Euclidean case, we obtain exponential decay for higher-order time and spatial derivatives of the solution in the damped cases, at least for large timeframes $t\geq 1$; see Theorem 2.7, 2.8, and 2.9. This is because of the discrete spectrum of the positive operator $\mathcal{L}$. However, such estimates are not possible in the case of operator having continuous spectrum; see [Bui, Iketha:struc].
  2. (2)
@@
    The absence of decay or the presence of polynomial-time decay in the interval $t\in(0,1]$ is generally inadequate to demonstrate global well-posedness. The estimates in Theorem 2.5 are best suited to prove global well-posedness, while the estimates in Theorem 2.4 can be used to obtain at least local well-posedness.
@@
The proofs for the effective, critical, and non-effective damping cases in Theorem 2.4 will be provided as remarks, following the proofs of Theorems 2.7, 2.8, and 2.9, respectively. The case of viscoelastic-type damping in Theorem 2.4 will be provided independently.
@@ 3. Preliminaries
The key components of Fourier analysis in relation to our indicated space and the self-adjoint, positive operator $\mathcal{L}$ will be reviewed in this section, see [DMT17, Ruz16, ruzhansky2017nonharmonic, MR&JPVR] for a thorough exposition. More precisely, this section defines the $\mathcal{L}$-Fourier transform, $\mathcal{L}$-convolution and the Sobolev spaces associated with the operator $\mathcal{L}$ and its properties. At the end, we will conclude this section by recalling some important inequalities.
@@
$|\phi(\xi)|\leq C_{\phi,N}\langle\xi\rangle^{-N}\quad\text{for all }\xi\in\mathcal{I},$
where $\langle\xi\rangle=(1+\lambda_{\xi})^{\frac{1}{2}}$. The topology on $\mathcal{S}(\mathcal{I})$ is given by the seminorms $p_{k},k\in\mathbb{N}_{0}$ and is defined by
$p_{k}(\phi):=\sup_{\xi\in\mathcal{I}}\langle\xi\rangle^{k}|\phi(\xi)|.$
@@
$\widehat{f}(\xi):=\left(\mathcal{F}_{\mathcal{L}}f\right)(\xi)=\langle f,u_{\xi}\rangle.$
@@
The Plancherel’s formula is given by
$\|f\|_{\mathcal{H}}=\left(\sum_{\xi\in\mathcal{I}}|\widehat{f}(\xi)|^{2}\right)^{1/2}.$ (3.1)
The $\mathcal{L}$-convolution of $f_{1},f_{2}\in H_{\mathcal{L}}^{\infty}$ is defined by
@@
We now recall Sobolev spaces generated by the operator $\mathcal{L}$. For $f\in\mathcal{D}^{\prime}(\mathcal{L})$ and $s>0,$ we say that $f\in\mathcal{H}_{\mathcal{L}}^{s}$ iff $\mathcal{L}^{s/2}f\in\mathcal{H}$ and the associated Sobolev norm is defined as
$\|f\|_{\mathcal{H}_{\mathcal{L}}^{s}}:=\|\mathcal{L}^{s/2}f\|_{\mathcal{H}}=\left(\sum_{\xi\in\mathcal{I}}\lambda_{\xi}^{s}|\widehat{f}(\xi)|^{2}\right)^{1/2}.$
@@
  1. (1)
    For $x\in[0,\frac{1}{4}]$, the following inequalities hold:
    $4x\leq-1+\sqrt{1-4x}\leq-2x\quad\text{and}\quad-2\leq-1-\sqrt{1-4x}\leq-1.$ (3.3)
  2. (2)
    For any $\gamma,\delta>0$, the function $f(t)=t^{\gamma}e^{-\delta t}$ and $f(t)=(1+t)^{\gamma}e^{-\delta t}$ is bounded on $0,\infty)$. Furthermore, for any $\gamma>0$ and $\beta>\delta>0$, we have
    for some positive constant $C(\gamma,\delta)>0$.
  3. (3)
    Since ensuring that $\delta$ remains independent of $\xi\in\mathcal{I}$ is adequate for our calculations, we will suppress the dependence of $\delta$ on the parameters $a$ and $\gamma$.
@@ 4. Preparatory Estimates
$\text{undamped i.e., }\theta=0:~~\tau_{\pm}$ $=\begin{cases}-\frac{1}{2}\pm\sqrt{\frac{1}{4}-\lambda_{\xi}^{\sigma}}&\text{if }\lambda_{\xi}^{\sigma}<\frac{1}{4},\\ -\frac{1}{2}&\text{if }\lambda_{\xi}^{\sigma}=\frac{1}{4},\\ -\frac{1}{2}\pm i\sqrt{\lambda_{\xi}^{\sigma}-\frac{1}{4}}&\text{if }\lambda_{\xi}^{\sigma}>\frac{1}{4},\end{cases}$ (4.2)
$\text{effective-damping i.e., }2\theta\in(0,\sigma):~~\tau_{\pm}$ $=\begin{cases}-\frac{1}{2}\lambda_{\xi}^{\theta}\pm\sqrt{\frac{1}{4}\lambda_{\xi}^{2\theta}-\lambda_{\xi}^{\sigma}}&\text{if }\lambda_{\xi}^{\sigma-2\theta}<\frac{1}{4},\\ -\frac{1}{2}\lambda_{\xi}^{\theta}&\text{if }\lambda_{\xi}^{\sigma-2\theta}=\frac{1}{4},\\ -\frac{1}{2}\lambda_{\xi}^{\theta}\pm i\sqrt{\lambda_{\xi}^{\sigma}-\frac{1}{4}\lambda_{\xi}^{2\theta}}&\text{if }\lambda_{\xi}^{\sigma-2\theta}>\frac{1}{4},\end{cases}$ (4.3)
$\text{critical damping i.e., }2\theta=\sigma:~~\tau_{\pm}$ $=-\frac{1}{2}\lambda_{\xi}^{\theta}\pm i\frac{\sqrt{3}}{2}\lambda_{\xi}^{\theta},$ (4.4)
$\text{non-effective-damping i.e., }2\theta\in(\sigma,2\sigma):~~\tau_{\pm}$ $=\begin{cases}-\frac{1}{2}\lambda_{\xi}^{\theta}\pm i\sqrt{\lambda_{\xi}^{\sigma}-\frac{1}{4}\lambda_{\xi}^{2\theta}}&\text{if }\lambda_{\xi}^{2\theta-\sigma}<4,\\ -\frac{1}{2}\lambda_{\xi}^{\theta}&\text{if }\lambda_{\xi}^{2\theta-\sigma}=4,\\ -\frac{1}{2}\lambda_{\xi}^{\theta}\pm\sqrt{\frac{1}{4}\lambda_{\xi}^{2\theta}-\lambda_{\xi}^{\sigma}}&\text{if }\lambda_{\xi}^{2\theta-\sigma}>4,\end{cases}$ (4.5)
$\text{viscoelastic type damping i.e., }\theta=\sigma:~~\tau_{\pm}$ $=\begin{cases}-\frac{1}{2}\lambda_{\xi}^{\theta}\pm\frac{i}{2}\sqrt{4\lambda_{\xi}^{\theta}-\lambda_{\xi}^{2\theta}}&\text{if }\lambda_{\xi}^{\theta}<4,\\ -\frac{1}{2}\lambda_{\xi}^{\theta}&\text{if }\lambda_{\xi}^{\theta}=4,\\ -\frac{1}{2}\lambda_{\xi}^{\theta}\pm\frac{1}{2}\sqrt{\lambda_{\xi}^{2\theta}-4\lambda_{\xi}^{\theta}}&\text{if }\lambda_{\xi}^{\theta}>4.\end{cases}$ (4.6)
@@
where $_{\mathcal{L}}$ is the convolution defined in (3.2), equivalently
@@
$\partial^{k}_{t}\widehat{E}_{1}(t,\xi)$ $=\begin{cases}\frac{(\tau_{+})^{k}e^{\tau_{+}t}-(\tau_{-})^{k}e^{\tau_{-}t}}{\tau_{+}-\tau_{-}}&\text{if }\tau_{+}\neq\tau_{-},\\ (\tau_{+})^{k-1}(k+t\tau_{+})e^{\tau_{+}t}&\text{if }\tau_{+}=\tau_{-}.\end{cases}$ (4.11)
@@ 5. Proof of main results
This section presents the proof of our main results, which is reliant upon the estimates of $\widehat{E}_{0}(t,\xi)$ and $\widehat{E}_{1}(t,\xi)$, as well as their spatial and temporal derivatives. The proofs of Theorem 2.4 and 2.5 will be presented gradually, accompanied by a set of remarks subsequent to the derivation of decay estimates in each respective case.
@@
where we have utilized the relation (3.3) to estimate the exponential factor and the fact that $\{\lambda_{\xi}\}_{\xi\in\mathcal{I}}$ is discrete, which allows us to bound $\frac{1}{\sqrt{\frac{1}{4}-\lambda_{\xi}^{\sigma}}}$ uniformly. Consequently, the representation (4.7) along with the above estimates gives
@@
where the constant $\delta$ is obtained by using the relation (3.5) which is due to discrete spectrum. Furthermore, for any $\beta>0$ we have
@@
since $\lambda_{\xi}^{\beta}$ is uniformly bounded on $\mathcal{R}_{2}$. Recalling the relation (4.11) for $\widehat{E}_{1}(t,\xi)$, we have
@@
for some modified $\delta>0$, where we have used the facts that $\mathcal{L}$ has a discrete spectrum with no cluster point, and $\frac{\left(-\frac{1}{2}+\sqrt{\frac{1}{4}-\lambda_{\xi}^{\sigma}}\right)^{k}}{2\sqrt{\frac{1}{4}-\lambda_{\xi}^{\sigma}}}$ and $\frac{\left(-\frac{1}{2}-\sqrt{\frac{1}{4}-\lambda_{\xi}^{\sigma}}\right)^{k}}{2\sqrt{\frac{1}{4}-\lambda_{\xi}^{\sigma}}}$ are uniformly bounded on $\mathcal{R}_{2}$. Consequently, using the relation (4.10), we obtain
@@
since $\lambda_{\xi}^{\sigma}$ is uniformly bounded on $\mathcal{R}_{2}$. Combining the above estimate with the representation (4.9), we get
@@
$|\lambda_{\xi}^{\beta}\widehat{E}_{0}(t,\xi)|\lesssim\lambda_{\xi}^{\beta}e^{-\frac{1}{2}t}\quad\text{and}\quad|\lambda_{\xi}^{\beta}\widehat{E}_{1}(t,\xi)|\lesssim\frac{\lambda_{\xi}^{\beta}}{\sqrt{\lambda_{\xi}^{\sigma}-\frac{1}{4}}}e^{-\frac{1}{2}t}\lesssim\lambda_{\xi}^{\beta-\frac{\sigma}{2}}e^{-\frac{1}{2}t},$
since $\frac{1}{\sqrt{1-\frac{1}{4}\lambda_{\xi}^{-\sigma}}}$ is uniformly bounded due to discrete spectrum. Consequently, using the above estimates with (4.7) we obtain
@@
  • The Remark 2.3 is now more understandable. Specifically, the estimate (5.15) shows that the regions $\mathcal{R}_{2}$, $\mathcal{R}_{3}$, and $\mathcal{R}_{4}$ experience exponential decay. Nevertheless, the possibility of decay in the total energy norm $\|u(t)\|_{\mathcal{H}}$ is obstructed by the lack of any decay in $\mathcal{R}_{1}$.
@@
$\widehat{E}_{0}(t,\xi)$ $=\begin{cases}1&\text{if }\xi\in\mathcal{R}_{1},\\ \left[\cosh\left(\sqrt{\frac{1}{4}\lambda_{\xi}^{2\theta}-\lambda_{\xi}^{\sigma}}t\right)+\frac{\lambda_{\xi}^{\theta}\sinh\left(\sqrt{\frac{1}{4}\lambda_{\xi}^{2\theta}-\lambda_{\xi}^{\sigma}}t\right)}{2\sqrt{\frac{1}{4}\lambda_{\xi}^{2\theta}-\lambda_{\xi}^{\sigma}}}\right]e^{-\frac{1}{2}\lambda_{\xi}^{\theta}t}&\text{if }\xi\in\mathcal{R}_{2},\\ \left[1+\frac{1}{2}\lambda_{\xi}^{\theta}t\right]e^{-\frac{1}{2}\lambda_{\xi}^{\theta}t}&\text{if }\xi\in\mathcal{R}_{3},\\ \left[\cos\left(\sqrt{\lambda_{\xi}^{\sigma}-\frac{1}{4}\lambda_{\xi}^{2\theta}}t\right)+\frac{\lambda_{\xi}^{\theta}\sin\left(\sqrt{\lambda_{\xi}^{\sigma}-\frac{1}{4}\lambda_{\xi}^{2\theta}}t\right)}{2\sqrt{\lambda_{\xi}^{\sigma}-\frac{1}{4}\lambda_{\xi}^{2\theta}}}\right]e^{-\frac{1}{2}\lambda_{\xi}^{\theta}t}&\text{if }\xi\in\mathcal{R}_{4},\end{cases}$ (5.18)
@@
$\widehat{E}_{1}(t,\xi)$ $=\begin{cases}t&\text{if }\xi\in\mathcal{R}_{1},\\ \frac{\sinh\left(\sqrt{\frac{1}{4}\lambda_{\xi}^{2\theta}-\lambda_{\xi}^{\sigma}}t\right)}{\sqrt{\frac{1}{4}\lambda_{\xi}^{2\theta}-\lambda_{\xi}^{\sigma}}}e^{-\frac{1}{2}\lambda_{\xi}^{\theta}t}&\text{if }\xi\in\mathcal{R}_{2},\\ te^{-\frac{1}{2}\lambda_{\xi}^{\theta}t}&\text{if }\xi\in\mathcal{R}_{3},\\ \frac{\sin\left(\sqrt{\lambda_{\xi}^{\sigma}-\frac{1}{4}\lambda_{\xi}^{2\theta}}t\right)}{\sqrt{\lambda_{\xi}^{\sigma}-\frac{1}{4}\lambda_{\xi}^{2\theta}}}e^{-\frac{1}{2}\lambda_{\xi}^{\theta}t}&\text{if }\xi\in\mathcal{R}_{4}.\end{cases}$ (5.19)